Solve the following systems of equations

2dx/dt+dy/dt-3x-y=t

dx/dt+dy/dt-4x-y=et






1
Expert's answer
2021-08-10T12:59:35-0400
"2x'+y'-3x-y=t"

"x'+y'-4x-y=e^t"




"y'=-2x'+3x+y+t"

"y'=-x'+4x+y+e^t"


"-2x'+3x+y+t=y'=-x'+4x+y+e^t"

"x'+x=t-e^t"

"\mu(t)=e^t"

"e^tx'+e^tx=e^tt-e^te^t"

"(e^tx)'=e^tt-e^{2t}"

Integrate


"\int d(e^tx)=\int(e^tt-e^{2t})dt"

"\int e^ttdt=te^t-e^t+C_3"

"e^tx=te^t-e^t-\dfrac{1}{2}e^{2t}+C_1"

"x(t)=t-1-\dfrac{1}{2}e^t+C_1e^{-t}"

"x'=1-\dfrac{1}{2}e^t-C_1e^{-t}"

"2-e^t-2C_1e^{-t}+y'-3t+3+\dfrac{3}{2}e^t-3C_1e^{-t}-y=t"

"y'-y=4t-5-\dfrac{1}{2}e^t+5C_1e^{-t}"

"\mu(t)=e^{-t}"

"e^{-t}y'-e^{-t}y=4te^{-t}-5e^{-t}-\dfrac{1}{2}+5C_1e^{-2t}"

"(e^{-t}y)'=4te^{-t}-5e^{-t}-\dfrac{1}{2}+5C_1e^{-2t}"

Integrate


"\int d(e^{-t}y)=\int(4te^{-t}-5e^{-t}-\dfrac{1}{2}+5C_1e^{-2t})dt"

"e^{-t}y=-4te^{-t}-4e^{-t}+5e^{-t}-\dfrac{1}{2}t-\dfrac{5}{2}C_1e^{-2t}+C_2"

"y(t)=-4t+1-\dfrac{1}{2}te^{t}-\dfrac{5}{2}C_1e^{-t}+C_2e^{t}"



"x(t)=t-1-\dfrac{1}{2}e^t+C_1e^{-t}"

"y(t)=-4t+1-\dfrac{1}{2}te^{t}-\dfrac{5}{2}C_1e^{-t}+C_2e^{t}"